OpenStudy (anonymous):

. a bus travel at 36 km/h and arrives in its destination half an hour late. If it travels at 42 km/h, it arrives at the same destination half an hour earlier. Find the journey distance.

OpenStudy (ash2326):

@mr.tobby Both the cases the distance is the same, Let time taken usually be t hours First case the bus arrives half an hour late. So time taken by it (t+0.5) hours Second case the bus arrives half an hour early. Time taken (t-0.5) hours Distance covered in first case \[D=36(t+0.5)\] Distance covered in second case \[D=42(t-0.5)\] Equate both these \[36(t+0.5)=42(t-0.5)\] Can you find t from this ? @mr.tobby

OpenStudy (anonymous):

t=6.5 ? then?

OpenStudy (ash2326):

yes, now find distance using any of the two equations.

OpenStudy (anonymous):

use D=42(t-0.5) and D = 42(t+0.5)

OpenStudy (ash2326):

Check your second equation :)

OpenStudy (anonymous):

wrong typed,:)

OpenStudy (ash2326):

Use any of the two both will give you the same answer :)